# Gaussian Copula Models for Nonignorable Missing Data Using Auxiliary Marginal Quantiles

### v1

@@ 2.2 Implied Additive Nonignorable Missingness Mechanism
The Gaussian copula with known margins implies a specific nonignorable missingness mechanism, $p(R_{ij}=1\mid\boldsymbol{y}^{obs},\boldsymbol{y}^{mis},\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha},\{F_{j}\}_{j=1}^{p})$. To show this, we first note that when $(Y,R)$ are distributed according (2)–(3), any subset of these variables also follows a Gaussian copula (Joe, , 2014). For the joint distribution of $(Y,R_{j})$, let the corresponding $(p+1)\times(p+1)$ copula correlation matrix be $\boldsymbol{C}_{j}^{*}$. This comprises the $p\times p$ sub-matrix $\boldsymbol{C}_{\boldsymbol{y}}$ of $\boldsymbol{C}_{\boldsymbol{\theta}}$ corresponding to the study variables concatenated with the $(p+1)\times 1$ column vector $(\boldsymbol{C}_{\boldsymbol{y}r_{j}},1)^{t}$ and $1\times(p+1)$ row vector $(\boldsymbol{C}_{r_{j}\boldsymbol{y}},1)$. Here, $\boldsymbol{C}_{\boldsymbol{y}r_{j}}$ comprises the entries of $\boldsymbol{C}_{\boldsymbol{\theta}}$ in the column for $R_{j}$, and $\boldsymbol{C}_{r_{j}\boldsymbol{y}}$ is its transpose. We have
$p(R_{ij}=1$ $\mid\boldsymbol{y}^{obs},\boldsymbol{y}^{mis},\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha},\{F_{j}\}_{j=1}^{p})=p(R_{ij}=1\mid\boldsymbol{y}_{i}^{obs},\boldsymbol{y}_{i}^{mis},\boldsymbol{C}_{j}^{*},\alpha_{r_{j}},\{F_{j}\}_{j=1}^{p})$ (7)
@@
Therefore, the marginal missingness mechanism is a probit regression on $(\boldsymbol{z}_{i}^{obs},\boldsymbol{z}_{i}^{mis})$.
The expression for $\alpha^{*}_{ij}$ reveals a connection to the additive nonignorable (AN) missingness mechanism. For any generic observation $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{ip})$ comprising observed and missing components $(\boldsymbol{x}_{i}^{obs},\boldsymbol{x}_{i}^{mis})$, the AN missingness mechanism holds for some $X_{j}$ when
@@
with $g$ satisfying $\lim_{\beta_{0}\rightarrow-\infty}g(\beta_{0})=0$ and $\lim_{\beta_{0}\rightarrow\infty}g(\beta_{0})=1$. Special cases of AN missingness mechanisms include an itemwise conditionally independent nonresponse (Sadinle and Reiter, , 2017) mechanism when $\beta_{j}=0$, and a MCAR mechanism when $\beta_{k}=0$ for all $k=1,\dots,p$. Common link functions $g$ include the logistic and probit (Hirano et al., , 2001).
Lemma 1 formally connects the model for $R_{ij}$ in (8) to the AN missingness mechanism in (10). Unlike the formulation in Hirano et al., (2001), the AN missingness mechanism for the Gaussian copula model has additivity on the latent scale.
@@
Of course, it is generally impossible to determine whether any specific missingness mechanism holds in practice (Molenberghs et al., , 2008). However, if additive nonignorability on the latent scale does not hold, e.g., the true model for $z_{r_{j}}$ includes interactions between $z_{j}$ and other latent variables, the Gaussian copula with fixed margins may not offer reliable inferences or sensitivity analyses. Developing methods and sensitivity analyses for nonignorable missingness that is not AN are topics for future research
3 Using Auxiliary Quantiles
@@
Figure 1: Using the set of auxiliary information $\mathcal{A}_{j}$ to create non-overlapping intervals $\{\mathcal{I}^{q}_{j}\}_{q=1}^{4}$ partitioning the support of $Y_{j}$. For $y_{ij}\in\mathcal{I}^{q}_{j}$, the auxiliary quantiles defining this interval pre-determine the interval containing $z_{ij}$ on the latent Gaussian scale. Here, $\mathcal{A}_{j}=\{F_{j}^{-1}(0)=-5,F_{j}^{-1}(0.25)=-2,F_{j}^{-1}(0.50)=0,F_{j}^{-1}(0.75)=3,F_{j}^{-1}(1)=5\}$.
@@ 3.1 Estimation of the Copula Correlation
The partial mapping between $y_{ij}^{obs}$ and $z_{ij}^{obs}$ in Figure 1 provides the basis for estimating the copula correlation. Using the intervals in Figure 1, define the binning function
@@
In what follows, we suppress the dependence of $b_{ij}^{obs}(y_{ij}^{obs})$ on $y_{ij}^{obs}$. Let $\boldsymbol{b}_{i}^{obs}=\{b_{ij}^{obs}:r_{ij}=0,j=1,\dots,p\}$ and $\boldsymbol{b}^{obs}=\{\boldsymbol{b}_{i}^{obs}:i=1,\dots,n\}$. The binning in (12) has the effect of coarsening the observed data (Heitjan and Rubin, , 1991; Miller and Dunson, , 2018) based on the intervals defined by $\mathcal{A}$. Thus, we require a likelihood for $(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$ using $(\boldsymbol{b}^{obs},\boldsymbol{r})$.
@@
The equivalence in (13) is by construction; observing $\boldsymbol{b}^{obs}$ implies that $\boldsymbol{z}^{obs}$ must belong to $\mathcal{D}(\boldsymbol{b}^{obs})$ conditional on $\mathcal{A}$. We refer to (13) as the extended quantile likelihood, abbreviated as EQL. The EQL also enables inclusion of discrete $Y_{j}$ in the copula model, as $\boldsymbol{b}^{obs}$ and $\mathcal{A}$ can be constructed from discrete support. Extensions of the EQL to incorporate unordered categorical study variables are discussed in the supplement.
@@
A similar estimation strategy for $\boldsymbol{C}_{\boldsymbol{\theta}}$ is employed under the extended rank (RL) and rank-probit (RPL) likelihoods (Hoff, , 2007; Feldman and Kowal, , 2022), which target posterior inference for the Gaussian copula correlation by conditioning on the set of latent variables consistent with the multivariate ranks on the observed variables. However, the EQL and the RL/RPL make different uses of their conditioning events. Under the EQL event, $\mathcal{A}$ partially locates $\boldsymbol{z}^{obs}$. By comparison, the RL/RPL event does not restrict where $\boldsymbol{z}^{obs}$ lies in latent space, as long as the orderings of the individual $z_{ij}^{obs}$ are consistent with the ranks of $y_{ij}^{obs}$. As a result, inferences for $\boldsymbol{C}_{\boldsymbol{\theta}}$ under AN missingness may be biased for the RL/RPL. We demonstrate this empirically in Section 4.1.
To estimate the model, we utilize a Gibbs sampler with data augmentation (Chib and Greenberg, , 1998), alternating sampling $(\boldsymbol{z}^{obs},\boldsymbol{z}^{mis},\boldsymbol{z}_{\boldsymbol{r}})\mid\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha}$ and $\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha}\mid\boldsymbol{(}\boldsymbol{z}^{obs},\boldsymbol{z}^{mis},\boldsymbol{z}_{\boldsymbol{r}})$. Algorithm 1 summarizes the two steps for an arbitrary specification of $p(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$. We present details for the $p(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$ used in the analyses in Section 4 and the supplement. In the algorithm, the subscript $(-j)$ in a vector indicates that vector without the $j$th element and in the column (row) index of a matrix indicates exclusion of the elements corresponding to the $j$th column (row) of that matrix. The subscript $-(jj)$ in a matrix indicates exclusion of all row and column elements for the $j$th variable in that matrix.
@@
  where $\mu_{ij}=\alpha_{j}+\boldsymbol{C}_{j(-j)}\boldsymbol{C}^{-1}_{-(jj)}(\boldsymbol{z}_{i(-j)}-\boldsymbol{\alpha}_{-j})$ and $\sigma_{j}^{2}=\boldsymbol{C}_{jj}-\boldsymbol{C}_{j(-j)}\boldsymbol{C}^{-1}_{-(jj)}\boldsymbol{C}_{(-j)j}$
@@
Algorithm 1 offers significant computational advantages over similar samplers for RL Gaussian copula models. Typically, $m_{j}<<n$, so the upper and lower truncation regions in Step 1 of Algorithm 1 are shared by many observations. Consequently, the data augmentation is computationally efficient: for all $b_{ij}^{obs}=q$, corresponding $z_{ij}^{obs}$ may be sampled simultaneously using truncated normal distributions. This enables reasonable computation times for (moderately) large $n$; for example, we fit the copula model to the North Carolina data comprising nearly 170,000 children. By contrast, the computational complexity of RL Gibbs samplers depends on the number of unique marginal ranks for each $Y_{j}$, which may approach $n$. Because of these computational benefits, it can be advantageous to specify auxiliary quantiles for $Y_{j}$ with no or MCAR missingness, for example, by letting $\mathcal{A}_{j}$ for those variables comprise a set of empirical quantiles. We note that empirical quantiles may be biased if the missingness mechanism is not MCAR. For such $Y_{j}$, analysts should specify a small set of auxiliary quantiles believed to closely approximate their corresponding true quantiles and conduct sensitivity analysis to alternative specifications of $\mathcal{A}_{j}$.
@@ 3.2 Imputation with Limited Auxiliary Information
The draws of (17) provide estimates of uncertainty about the intermediate quantiles. When few individuals have $b_{ij}^{obs}=q$, which may occur for intervals constructed at quantiles in the tails, we expect higher uncertainty in the draws of $\tilde{F}_{j}(y_{j}^{q})$. This is borne out in the simulations of Section 4. Algorithm 2 summarizes the process of estimating $\{\tilde{F}_{j}(y_{j}^{q})\}$.
@@
Using (17) in addition to $\mathcal{A}_{j}$ to approximate $F_{j}$ has advantages over approximating $F_{j}$ based on the quantiles in $\mathcal{A}_{j}$ alone, as done in the EQL. First, $\{y_{j}^{q}\}$ lends more information to the discrete approximation, helping the estimator better capture features of each marginal. Second, the draws of $\Phi(Z_{j}^{q})$ in (17) propagate uncertainty about ${F}_{j}(y_{j}^{q})$ to the imputations, whereas the EQL interpolation of ${F}_{j}(y_{j}^{q})$ is deterministic. For these reasons, we recommend employing the EHQL for copula estimation and Algorithm 2 for marginal CDF estimation whenever imputation is needed. We note that Feldman and Kowal, (2024) used a strategy similar to (17) for the RL Gaussian mixture copula under MAR mechanisms.
@@ 4 Simulation Studies
In this section, we present results of simulation studies evaluating (i) the impact of the level of detail in the marginal distributions on the quality of inferences and (ii) the repeated sampling performance of the model as a multiple imputation engine compared to alternative methods that do not use the auxiliary information for imputations.
@@
Here, $\boldsymbol{\Lambda}$ is a $(2p)\times k$ matrix of factor loadings possibly with $k<<2p$; $\boldsymbol{\eta}_{i}$ is a $k\times 1$ vector of factors; and, $\boldsymbol{\Sigma}=diag(\sigma_{1}^{2},\dots,\sigma_{2p}^{2})$. By specifying $\boldsymbol{\eta}_{i}\sim N_{k}(\boldsymbol{0},\boldsymbol{I}_{k})$, where $\boldsymbol{I}_{k}$ is the $k\times k$ identity matrix, marginally we have $\boldsymbol{z}_{i}\sim N_{2p}(\boldsymbol{\alpha},\boldsymbol{\Omega})$, where $\boldsymbol{\Omega}=\boldsymbol{\Lambda}\boldsymbol{\Lambda}^{\intercal}+\boldsymbol{\Sigma}$ is the reduced rank covariance. Thus, $\boldsymbol{\theta}=(\boldsymbol{\Lambda},\boldsymbol{\Sigma})$. Given posterior samples of $\boldsymbol{\Omega}$, samples of $\boldsymbol{C}_{\boldsymbol{\theta}}$ are obtained by scaling $\boldsymbol{\Omega}$ into correlations. The prior on the factor loadings provides shrinkage, automating rank selection (Bhattacharya and Dunson, , 2011). In addition, the components of $\boldsymbol{z}_{i}$ are independent conditional on $\boldsymbol{\eta}_{i}$, which benefits computation, especially in Step 1 of Algorithm 1. The full hierarchical specification of (18) is available in the supplement.

### v2

@@ 2.2 Implied Additive Nonignorable Missingness Mechanism
The Gaussian copula with known margins implies a specific nonignorable missingness mechanism, $p(R_{ij}=1\mid\boldsymbol{y}^{obs},\boldsymbol{y}^{mis},\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha},\{F_{j}\}_{j=1}^{p})$. To show this, we first note that when $(Y,R)$ are distributed according (2)–(3), any subset of these variables also follows a Gaussian copula (Joe, 2014). For the joint distribution of $(Y,R_{j})$, let the corresponding $(p+1)\times(p+1)$ copula correlation matrix be $\boldsymbol{C}_{j}^{*}$. This comprises the $p\times p$ sub-matrix $\boldsymbol{C}_{\boldsymbol{y}}$ of $\boldsymbol{C}_{\boldsymbol{\theta}}$ corresponding to the study variables concatenated with the $(p+1)\times 1$ column vector $(\boldsymbol{C}_{\boldsymbol{y}r_{j}},1)^{t}$ and $1\times(p+1)$ row vector $(\boldsymbol{C}_{r_{j}\boldsymbol{y}},1)$. Here, $\boldsymbol{C}_{\boldsymbol{y}r_{j}}$ comprises the entries of $\boldsymbol{C}_{\boldsymbol{\theta}}$ in the column for $R_{j}$, and $\boldsymbol{C}_{r_{j}\boldsymbol{y}}$ is its transpose. We have
$p(R_{ij}=1$ $\mid\boldsymbol{y}^{obs},\boldsymbol{y}^{mis},\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha},\{F_{j}\}_{j=1}^{p})=p(R_{ij}=1\mid\boldsymbol{y}_{i}^{obs},\boldsymbol{y}_{i}^{mis},\boldsymbol{C}_{j}^{*},\alpha_{r_{j}},\{F_{j}\}_{j=1}^{p})$
@@
Therefore, the marginal missingness mechanism is a probit regression on $(\boldsymbol{z}_{i}^{obs},\boldsymbol{z}_{i}^{mis})$ with parameters expressed as a function of the copula correlation.
The expression for $\alpha^{*}_{ij}$ reveals a connection to the AN missingness mechanism. For any generic observation $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{ip})$ comprising observed and missing components $(\boldsymbol{x}_{i}^{obs},\boldsymbol{x}_{i}^{mis})$, the AN missingness mechanism holds for some $X_{j}$ when
@@
with $g$ satisfying $\lim_{\beta_{0}\rightarrow-\infty}g(\beta_{0})=0$ and $\lim_{\beta_{0}\rightarrow\infty}g(\beta_{0})=1$. The key feature of AN mechanisms is the additivity in $g$; that is, (9) does not include interactions between $X_{j}$ and other study variables. This additivity is required for model identification (Hirano et al., 2001; Sadinle and Reiter, 2019). Special cases of AN mechanisms include itemwise conditionally independent nonresponse (Sadinle and Reiter, 2017) when $\beta_{j}=0$ and MCAR when $\beta_{k}=0$ for all $k=1,\dots,p$. Common link functions $g$ include the logistic and probit (Hirano et al., 2001).
Lemma 1 formally connects the model for $R_{ij}$ in (7) to the AN missingness mechanism in (9). Unlike the formulation in Hirano et al. (2001), the AN missingness mechanism for the Gaussian copula model has additivity on the latent scale.
@@
Of course, it is generally impossible to determine whether any specific missingness mechanism holds in practice (Molenberghs et al., 2008). However, if additive nonignorability on the latent scale does not hold, e.g., the true model for $z_{r_{j}}$ includes interactions between $z_{j}$ and other latent variables, the Gaussian copula with fixed margins may not offer reliable inferences or sensitivity analyses. Developing methods for sensitivity analyses with other nonignorable missingness mechanisms is a topic for future research.
3 Leveraging Incomplete Auxiliary Quantiles
@@
Figure 1: Using auxiliary information $\mathcal{A}_{j}$ to create non-overlapping intervals $\{\mathcal{I}^{q}_{j}\}_{q=1}^{4}$ partitioning the support of $Y_{j}$. For $y_{ij}\in\mathcal{I}^{q}_{j}$, the auxiliary quantiles defining this interval pre-determine the interval containing $z_{ij}$ on the latent Gaussian scale. Here, $\mathcal{A}_{j}=\{F_{j}^{-1}(0)=-5,F_{j}^{-1}(0.25)=-2,F_{j}^{-1}(0.50)=0,F_{j}^{-1}(0.75)=3,F_{j}^{-1}(1)=5\}$.
@@ 3.1 Estimation of the Copula Correlation
The partial mapping between $y_{ij}^{obs}$ and $z_{ij}^{obs}$ in Figure 1 provides the basis for estimating $\boldsymbol{C}_{\boldsymbol{\theta}}$. Using the intervals in Figure 1, define the binning function
@@
In what follows, we suppress the dependence of $b_{ij}^{obs}(y_{ij}^{obs})$ on $y_{ij}^{obs}$. Let $\boldsymbol{b}_{i}^{obs}=\{b_{ij}^{obs}:r_{ij}=0,j=1,\dots,p\}$ and $\boldsymbol{b}^{obs}=\{\boldsymbol{b}_{i}^{obs}:i=1,\dots,n\}$. The binning in (11) has the effect of coarsening the observed data (Heitjan and Rubin, 1991; Miller and Dunson, 2018) based on the intervals defined by $\mathcal{A}$. Thus, we require a likelihood for $(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$ using $(\boldsymbol{b}^{obs},\boldsymbol{r})$.
@@
The equivalence in (12) is by construction; observing $\boldsymbol{b}^{obs}$ implies that $\boldsymbol{z}^{obs}$ must belong to $\mathcal{D}(\boldsymbol{b}^{obs})$ conditional on $\mathcal{A}$. We refer to (12) as the extended quantile likelihood, abbreviated as EQL. The EQL also enables inclusion of discrete $Y_{j}$ in the copula model, as $\boldsymbol{b}^{obs}$ and $\mathcal{A}$ can be constructed from discrete support. We discuss including unordered categorical variables in the EQL in the supplement, Section D.
@@
The EQL resembles the extended rank (RL) and rank-probit (RPL) likelihoods (Hoff, 2007; Feldman and Kowal, 2022), which target posterior inference for the Gaussian copula correlation by conditioning on the set of latent variables consistent with the multivariate ranks of the observed data values. However, the EQL and the RL/RPL make different uses of their conditioning events. Under the EQL event, $\mathcal{A}$ partially locates $\boldsymbol{z}^{obs}$. By comparison, the RL/RPL event does not restrict where $\boldsymbol{z}^{obs}$ lies in latent space, as long as the orderings of the individual $z_{ij}^{obs}$ are consistent with the ranks of $y_{ij}^{obs}$. As a result, inferences for $\boldsymbol{C}_{\boldsymbol{\theta}}$ under AN missingness may be biased for the RL/RPL, as we demonstrate in Section 4.1..
To estimate the model, we utilize a Gibbs sampler with data augmentation (Chib and Greenberg, 1998), alternating sampling $(\boldsymbol{z}^{obs},\boldsymbol{z}^{mis},\boldsymbol{z}_{\boldsymbol{r}})\mid\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha}$ and $\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha}\mid\boldsymbol{(}\boldsymbol{z}^{obs},\boldsymbol{z}^{mis},\boldsymbol{z}_{\boldsymbol{r}})$. Algorithm 1 summarizes the two steps for an arbitrary specification of $p(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$. We present details for the $p(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$ used in the analyses in Section 4 and Section E of the supplement. In the algorithm, the subscript $(-j)$ in a vector indicates that vector without the $j$th element and in the column (row) index of a matrix indicates exclusion of the elements corresponding to the $j$th column (row) of that matrix. The subscript $-(jj)$ in a matrix indicates exclusion of all row and column elements for the $j$th variable in that matrix.
@@
  where $\mu_{ij}=\alpha_{j}+\boldsymbol{C}_{j(-j)}\boldsymbol{C}^{-1}_{-(jj)}(\boldsymbol{z}_{i(-j)}-\boldsymbol{\alpha}_{-j})$, $\sigma_{j}^{2}=\boldsymbol{C}_{jj}-\boldsymbol{C}_{j(-j)}\boldsymbol{C}^{-1}_{-(jj)}\boldsymbol{C}_{(-j)j}$
@@
Algorithm 1 offers significant computational advantages over similar samplers for RL/RPL Gaussian copula models. Typically, $m_{j}<<n$, so the upper and lower truncation regions in Step 1 of Algorithm 1 are shared by many observations. Consequently, the data augmentation is computationally efficient: for all $b_{ij}^{obs}=q$, corresponding $z_{ij}^{obs}$ may be sampled simultaneously using truncated normal distributions. This enables reasonable computation times for (moderately) large $n$; for example, we fit the copula model to the North Carolina data comprising nearly 170,000 children. By contrast, the computational complexity of RL Gibbs samplers depends on the number of unique marginal ranks for each $Y_{j}$, which may approach $n$. Because of these computational benefits, it can be advantageous to specify auxiliary quantiles for $Y_{j}$ with no or MCAR missingness, for example, by letting $\mathcal{A}_{j}$ for those variables comprise a set of empirical quantiles. We note that empirical quantiles may be biased if the missingness mechanism is not MCAR. For such $Y_{j}$, analysts should specify a small set of auxiliary quantiles believed to closely approximate their corresponding true quantiles and conduct sensitivity analysis to alternative specifications of $\mathcal{A}_{j}$.
@@ 3.2 Imputation with Limited Auxiliary Information
The draws of (16) provide estimates of uncertainty in the intermediate quantiles. When few individuals have $b_{ij}^{obs}=q$, which may occur for intervals constructed at quantiles in the tails, we expect higher uncertainty in the draws of $\tilde{F}_{j}(y_{j}^{q})$. This is borne out in the simulations of Section 4. Algorithm 2 summarizes the process for estimating the $\{\tilde{F}_{j}(y_{j}^{q})\}$.
@@
Using (16) in addition to $\mathcal{A}_{j}$ to approximate $F_{j}$ has advantages over approximating $F_{j}$ based on the quantiles in $\mathcal{A}_{j}$ alone, as done in the EQL. First, $\{y_{j}^{q}\}$ lends more information to the discrete approximation, helping the estimator better capture features of each marginal. Second, the draws of $\Phi(Z_{j}^{q})$ in (16) propagate uncertainty about ${F}_{j}(y_{j}^{q})$ to the imputations, whereas the EQL interpolation of ${F}_{j}(y_{j}^{q})$ is deterministic. For these reasons, we recommend employing the EHQL for copula estimation and Algorithm 2 for marginal CDF estimation whenever imputation is needed. We note that Feldman and Kowal (2024) use a strategy akin to what we do in (16) for the RL Gaussian mixture copula under MAR mechanisms. However, their method cannot leverage auxiliary marginal information and hence does not address the nonignorable missingness setting we consider here.
@@ 4 Simulation Studies
In this section, we present results of simulation studies evaluating (i) the impact of the level of detail in the marginal distributions on the quality of inferences and (ii) the repeated sampling performance of the model as a multiple imputation engine compared to alternative methods that do not use the auxiliary information.
@@
Here, $\boldsymbol{\Lambda}$ is a $(2p)\times k$ matrix of factor loadings possibly with $k<<2p$; $\boldsymbol{\eta}_{i}$ is a $k\times 1$ vector of factors; and, $\boldsymbol{\Sigma}=diag(\sigma_{1}^{2},\dots,\sigma_{2p}^{2})$. By specifying $\boldsymbol{\eta}_{i}\sim N_{k}(\boldsymbol{0},\boldsymbol{I}_{k})$, where $\boldsymbol{I}_{k}$ is the $k\times k$ identity matrix, marginally we have $\boldsymbol{z}_{i}\sim N_{2p}(\boldsymbol{\alpha},\boldsymbol{\Omega})$, where $\boldsymbol{\Omega}=\boldsymbol{\Lambda}\boldsymbol{\Lambda}^{\intercal}+\boldsymbol{\Sigma}$ is the reduced rank covariance. Thus, $\boldsymbol{\theta}=(\boldsymbol{\Lambda},\boldsymbol{\Sigma})$. Given posterior samples of $\boldsymbol{\Omega}$, samples of $\boldsymbol{C}_{\boldsymbol{\theta}}$ are obtained by scaling $\boldsymbol{\Omega}$ into correlations. The prior on the factor loadings provides shrinkage, automating rank selection (Bhattacharya and Dunson, 2011). In addition, the components of $\boldsymbol{z}_{i}$ are independent conditional on $\boldsymbol{\eta}_{i}$, which benefits computation, especially in Step 1 of Algorithm 1. The full hierarchical specification of (17) is available in Section E of the supplement.